# MIBoost: A Gradient Boosting Algorithm for Variable Selection After Multiple Imputation

### v4

@@ 6 Conclusion
Furthermore, the miselect package allows for weighting observations to account for the fact that stacked datasets consist of $M\times n$ observations. For both SaLASSO and SaENET, we chose observation weights of $1/M$, as recommended in [4]. An alternative weighting approach—scaling each observation by the proportion of missing values across all variables—was also discussed in [4]. However, the authors acknowledged that this method would shift the analysis closer to a complete‑case analysis, which runs counter to the purpose of multiple imputation. We therefore did not adopt this approach.
Finally, since we did not apply standard LASSO or elastic nets but rather their adaptive versions, we needed to define adaptive weight vectors for each run of SaLASSO or SaENET. To do so, we first fitted a preliminary stacked elastic net with equal weights (referred to as SENET in Du et al. [4]). We then extracted the estimated coefficients of all candidate covariates and used them to compute the corresponding adaptive weight vectors.

### v5

@@ C Specifications for SaLASSO and SaENET
Furthermore, the miselect package allows for weighting observations to account for the fact that stacked datasets consist of $M\times n$ observations. For both SaLASSO and SaENET, we chose observation weights of $1/M$, as recommended in [undefg]. An alternative weighting approach—scaling each observation by the proportion of missing values across all variables—was also discussed in [undefg]. However, the authors acknowledged that this method would shift the analysis closer to a complete-case analysis, which runs counter to the purpose of multiple imputation. We therefore did not adopt this approach.
Finally, since we did not apply standard LASSO or elastic nets but rather their adaptive versions, we needed to define adaptive weight vectors for each run of SaLASSO or SaENET. To do so, we first fitted a preliminary stacked elastic net with equal weights (referred to as SENET in [undefg]). We then extracted the estimated coefficients of all candidate covariates and used them to compute the corresponding adaptive weight vectors.